OpenStudy (anonymous):

Each morning Henry takes the subway from Town Center to Maple Station. As the map shows below Maple Station is 3 miles west and 5 miles north of Town Center. Find the total distance Henry travels from Town Center to Maple Station within 2 decimal places.

OpenStudy (anonymous):

so you would have to use pythagorean theorem, as you have two sides of the triangle 3 and 5. so you do 3^2 + 5^2 = 9 + 25 = 34 so the distance is the sqrt of 34

OpenStudy (anonymous):

5.83095 approximately